# Multimodal Fusion and Vision-Language Models: A Survey for Robot Vision

### v1

@@ 2.3.2 Visual Localization
Visual Localization plays a pivotal role in robot vision, enabling autonomous agents to accurately determine their position in complex and dynamic environments. Recent advancements in multimodal fusion, particularly deep-learning-based approaches, have significantly enhanced localization accuracy and robustness. These methods leverage complementary sensory inputs such as RGB images, depth data, and inertial measurements to address challenges like sensor noise, dynamic scenes, and varying lighting conditions [129, 130, 131]. Model,s like DeepVO [20] and D3VO [132] exemplify the integration of convolutional and recurrent neural networks, enabling end-to-end estimation of ego-motion and depth. Supervised methods excel in constrained settings by utilizing labeled datasets, while self-supervised approaches leverage geometric and photometric consistency to adapt effectively to unseen environments [133].
Hybrid frameworks combining classical geometric models with deep neural networks offer a promising solution to challenges like scale ambiguity and drift. By integrating learned depth and optical flow predictions [134], these systems achieve consistent pose estimations, improving monocular system accuracy. For example, D3VO incorporates uncertainty estimation into the visual odometry pipeline for robust performance in dynamic and occluded environments. This highlights the value of merging data-driven and model-based methods for reliable localization. Recent advances in scene coordinate regression and neural implicit representations are transforming the field. Scene coordinate regression techniques eliminate the need for explicit 3D maps by learning direct transformations from image pixels to 3D coordinates, simplifying localization. Neural implicit representations, like NeRF [135] and Semantic-NeRF [136], encode scene geometry and semantics into compact neural fields, enabling robust localization in new environments and showcasing the potential of multimodal fusion for bridging perception and mapping. Despite progress, challenges remain, such as the need for large-scale datasets, high computational costs, and limited generalization to real-world conditions. Future research will likely focus on improving scalability, real-time capabilities, and developing unified frameworks that integrate localization with high-level perception and decision-making tasks [137, 138].
@@ 2.5.1 Vision-Language-Action Models
Multimodal feature fusion is the core technology enabling the seamless transition from perception to action in robot manipulation tasks. Vision-Language-Action (VLA) models, which integrate visual perception, language understanding, and action planning, offer an efficient framework for addressing the challenges of complex manipulation tasks. The key innovation in VLA models lies in their ability to extract complementary features from different modalities and integrate them into a unified representation, enabling precise action planning and execution.
The visual modality is crucial for manipulation tasks, providing information about object position, shape, and layout. However, relying only on vision limits the robot’s task understanding. To address this, VLA models integrate language to process natural language instructions, bridging gaps in visual perception. For example, the RT-2 [152] model aligns visual and language representations via pretraining, allowing the robot to generate control commands from linguistic input using a Transformer architecture to map high-level semantics to actions. Additionally, the RoboMamba [153] model incorporates action dynamics into the multimodal fusion framework, optimizing vision, language, and action for task planning. The 3D-VLA [154] model enhances visual representation by generating 3D point clouds and depth maps, improving task precision and robustness with language-guided action generation in dynamic environments. Efficient alignment of multimodal features is critical, with OpenVLA [19] using Low-Rank Adaptation (LoRA) for parameter-efficient fusion, while DeeR-VLA [155] optimizes reasoning with a dynamic early-exit mechanism to reduce computational overhead. These methods enhance task success rates and enable efficient deployment on resource-constrained platforms. Furthermore, VoxPoser [156] introduces a language-driven affordance inference mechanism to generate 3D value maps for complex manipulation tasks, dynamically adjusting manipulation strategies through semantic reasoning for real-time adaptation in dynamic environments.
@@ 3.2 Cross-Modal Alignment and Representation Learning
Cross-modal alignment and representation learning are crucial for vision-language models, enabling different modalities (e.g., vision, text, audio) to interact in a shared embedding space. Alignment helps models map relationships between modalities, such as images to text or images to audio. Representation learning extracts useful features from each modality for various tasks. This combination improves performance in tasks like image captioning, visual question answering, and speech recognition, advancing multimodal capabilities. Cross-modal alignment maps information from modalities, such as images and text, into a shared space to understand their relationships [177]. However, challenges arise from differences in structure, representation, and data features between modalities. Images are high-dimensional, while text consists of discrete symbols, making alignment difficult. Other challenges include scale differences and noise in large-scale unlabeled data, which can affect alignment accuracy [178].
Common methods for representation learning include contrastive learning, self-supervised learning, and cross-modal generative models. Contrastive learning aims to minimize the distance between positive samples (same-category) and maximize it for negative samples (different-category). For cross-modal tasks, positive samples are visual and language pairs from the same context, while negative samples come from different contexts. The model is trained using a contrastive loss function to bring related modality pairs, like images and descriptions, closer in a shared embedding space while distancing unrelated pairs [179, 180]. The popular method CLIP [181] trains both vision and language models using contrastive learning, aligning images and text in a shared space for consistency. ALIGN [182] follows a similar approach, training cross-modal image and text data to enhance matching performance.
@@ 4.1 Scene Understanding Datasets
In the domain of autonomous driving, nuScenes [23] and the Waymo [236] Open Dataset are two iconic datasets. nuScenes [23] offers data from six cameras, five radars, and one LiDAR, covering 360° views and supporting 3D detection and tracking tasks across 23 categories. The Waymo [236] dataset leads in terms of geographic coverage and data scale, containing 12 million 3D bounding box annotations and supporting cross-domain generalization and multimodal sensor fusion research. Additionally, SemanticKITTI [24] focuses on semantic segmentation of LiDAR point clouds and provides an important benchmark for sequence-level semantic scene completion. The Cityscapes [237] dataset, with its high-resolution pixel-level annotations and complex urban street scene data, has become a crucial resource for urban semantic understanding.
For indoor scene understanding, datasets such as Matterport3D[26], ScanNet [238], NYU Depth V2 [239], and Hypersim [240] have been pivotal. Matterport3D [26] provides 10,800 panoramic RGB-D views, along with accurate 3D reconstructions and semantic annotations, making it widely applicable for semantic segmentation, 3D reconstruction, and multi-task learning. ScanNet [238], through its 2.5 million RGB-D frames across 1,513 scenes, utilizes crowdsourced annotations to provide a large-scale, high-quality dataset for semantic segmentation and 3D modeling. NYU Depth V2 [239], as one of the earliest RGB-D datasets, offers 1,449 images from 464 diverse indoor scenes with detailed semantic segmentation and object relationship annotations, making it instrumental for research on physical support reasoning and cluttered environments. In addition, the Hypersim dataset [240], utilizing high-fidelity synthetic techniques, generates 77,400 indoor scene images with detailed pixel-level annotations, light decomposition, and surface normals, serving as an important tool for simulation-to-reality (Sim-to-Real) transfer research.
@@
Datasets Scene Multimodal Data Venue Year
360+x [241] Indoor, Outdoor Video/Audio CVPR 2024
ScanQA [242] Indoor RGB/Text CVPR 2022
Hypersim [240] Indoor RGB/Depth ICCV 2021
NuScenes [23] Urban street RGB/Lidar/Radar CVPR 2020
Waymo [236] Outdoor RGB/Lidar CVPR 2020
Semantickitti [24] Urban street RGB/Lidar ICCV 2019
Matterport3D [26] Indoor RGB/Depth arxiv 2017
ScanNet [238] Indoor RGB/Depth CVPR 2017
Cityscapes [237] Urban street RGB/Depth CVPR 2016
NYUDv2 [239] Indoor RGB/Depth ECCV 2012
@@ 4.2 Robot Manipulation Datasets
In the domain of multimodal perception, Touch and Go [245] and VisGel [246] have pioneered research into the integration of tactile and visual data. Touch and Go [245] collects paired visual and tactile data from real-world environments and explores cross-modal relationships using self-supervised learning methods. VisGel [246], leveraging GelSight sensors, constructs bidirectional generation tasks between tactile and visual modalities, further validating the feasibility of inter-modal information transformation. These datasets provide valuable experimental resources for enhancing multimodal perception research. However, compared to ObjectFolder 2.0 [247], these tactile datasets are relatively smaller in scale. ObjectFolder 2.0 [247], with its high-quality virtual object data and implicit neural representation technology, achieves efficient rendering of visual, auditory, and tactile modalities, supporting virtual-to-reality transfer learning tasks. This makes it a significant breakthrough in the field of multimodal perception learning.
Overall, these datasets demonstrate unique strengths in modality design and task focus. DROID [25] and RT-1 [244] emphasize broad coverage across multi-task and multi-scene scenarios, while Open X-Embodiment provides a unified platform for multi-robot learning through task-agnostic data collection. Among the grasping task datasets, R2SGrasp [243] highlights domain adaptation for real-world scenarios, whereas Grasp-Anything-6D and Grasp-Anything++ [249] significantly improve grasping capabilities in complex environments by integrating language modalities and innovative model architectures. Tactile datasets like Touch and Go and VisGel offer unique perspectives on the integration of visual and tactile modalities, while ObjectFolder 2.0 [247] provides broader multimodal support and efficient virtual object rendering, laying the foundation for virtual-to-reality migration in perception tasks.
The complementarity of these datasets provides rich resources for multimodal robotic research. In the future, integrating task-agnostic multimodal platforms (e.g., Open X-Embodiment [250]) with task-specific fine-grained datasets (e.g., Grasp-Anything++ [249]) could help build more versatile and intelligent robotic systems, achieving comprehensive advancements from perception to decision-making.
@@ 5.1 Evaluation Metrics
where: - $mAP$ is the mean Average Precision, - $E_{i}$ represents the five error metrics: - mean Average Translation Error (mATE), - mean Average Scale Error (mASE), - mean Average Orientation Error (mAOE), - mean Average Velocity Error (mAVE), - mean Average Attribute Error (mAAE). These metrics ensure accurate assessment of multimodal fusion models’ capability to detect and localize objects.
@@
$GP=1-\frac{\text{Final distance to goal}}{\text{Initial distance to goal}}$ (12)
@@ 5.2 Performance comparison on 3D Object Detection
Methods Modality mAP mATE mASE mAOE mAVE NDS Car Truck C.V. Bus T.L. B.R. M.T. Bike Ped. T.C. VoxelNeXt [254] L 64.5 0.268 0.238 0.377 0.219 70.0 84.6 53.0 28.7 64.7 55.8 74.6 73.2 45.7 85.8 79.0 TransFusion-L [13] L 65.5 0.256 0.240 0.351 0.278 70.2 86.2 56.7 28.2 66.3 58.8 78.2 68.3 44.2 86.1 82.0 FocalFormer3D [255] L 68.7 0.254 0.242 0.340 0.218 72.6 87.2 57.0 34.4 69.6 64.9 77.8 76.2 49.6 88.2 82.3 UniM2AE-v2 [256] C 63.4 0.401 0.235 0.236 0.154 70.1 77.6 50.7 40.1 46.3 56.7 75.7 72.4 58.9 70.6 84.5 SparseBEV [257] C 60.3 0.425 0.239 0.311 0.172 67.5 76.3 49.2 35.6 44.2 53.4 76.8 64.0 53.7 65.8 83.7 RayDN [258] C 63.1 0.437 0.235 0.283 0.220 68.6 77.1 54.4 35.9 53.7 56.8 76.9 67.5 57.6 69.4 81.8 FusionPainting [83] C+L 68.1 0.256 0.236 0.346 0.274 71.6 87.1 60.8 30.0 68.5 61.7 71.8 74.7 53.5 88.3 85.0 AutoAlignV2 [259] C+L 68.4 0.245 0.233 0.311 0.258 72.4 87.0 59.0 33.1 69.3 59.3 78.0 72.9 52.1 87.6 85.1 IS-Fusion [90] C+L 76.5 0.230 0.231 0.267 0.223 77.4 89.8 67.8 44.5 77.6 68.3 81.8 85.3 65.6 93.4 91.1 UVTR-Multimodalit [84] C+L 67.1 0.306 0.245 0.351 0.225 71.1 87.5 56.0 33.8 67.5 59.5 73.0 73.4 54.8 86.3 79.6 3D-CVF [81] C+L 52.7 0.300 0.245 0.458 0.279 62.3 83.0 45.0 15.9 48.8 49.6 65.9 51.2 30.4 74.2 62.9 CMT [87] C+L 70.4 0.299 0.241 0.323 0.240 73.0 87.2 61.5 37.5 72.4 62.8 74.7 79.4 58.3 86.9 83.2 TransFusion [13] C+L 68.9 0.259 0.243 0.359 0.288 71.7 87.1 60.0 33.1 68.3 60.8 78.1 73.6 52.9 88.4 86.7 BEVFusion [88] C+L 70.2 0.261 0.239 0.329 0.260 72.9 88.6 60.1 39.3 69.8 63.8 80.0 74.1 51.0 89.2 86.5 DeepInteraction-e [86] C+L 75.6 0.235 0.233 0.328 0.226 76.3 88.3 64.3 44.7 74.2 66.0 83.5 85.4 66.4 92.5 90.9 SimpleBEV [260] C+L 75.7 0.236 0.235 0.283 0.143 77.6 89.0 69.2 75.2 70.5 44.4 91.7 81.2 69.1 87.9 78.4 DAL [261] C+L 72.0 0.253 0.238 0.334 0.174 74.8 89.1 60.2 73.3 65.8 34.6 89.6 81.7 58.5 86.6 80.6 MV2DFusion-e [262] C+L 77.9 0.237 0.225 0.247 0.192 78.8 91.1 69.7 45.7 76.7 70.4 83.0 87.1 72.2 93.3 90.3 RCBEVDet++ [263] C+R 67.3 0.341 0.234 0.241 0.147 72.7 83.5 56.0 45.0 54.1 57.9 76.9 78.4 60.2 76.4 85.0 CRN [101] C+R 57.5 0.416 0.264 0.456 0.365 62.4 78.4 52.6 29.2 57.0 48.3 64.7 63.4 47.8 58.2 74.9 RCM-Fusion [103] C+R 50.6 0.465 0.254 0.384 0.438 58.7 73.9 45.6 24.3 51.6 42.1 62.9 53.5 34.9 50.0 67.6 HyDRa [16] C+R 57.4 0.398 0.252 0.424 0.250 64.2 79.0 50.0 42.9 48.5 29.4 65.8 64.0 45.3 75.9 72.9 CenterPoint v2 [264] C+L+R 67.1 0.249 0.236 0.350 0.250 71.4 87.0 57.3 28.8 69.3 60.4 71.0 71.3 49.0 90.4 86.8

### v3

@@ 2.3.2 Visual Localization
Visual Localization plays a pivotal role in robot vision, enabling autonomous agents to accurately determine their position in complex and dynamic environments. Recent advancements in multimodal fusion, particularly deep-learning-based approaches, have significantly enhanced localization accuracy and robustness [134]. These methods leverage complementary sensory inputs such as RGB images, depth data, and inertial measurements to address challenges like sensor noise, dynamic scenes, and varying lighting conditions [135, 136, 137]. Model,s like DeepVO [20] and D3VO [138] exemplify the integration of convolutional and recurrent neural networks, enabling end-to-end estimation of ego-motion and depth. Supervised methods excel in constrained settings by utilizing labeled datasets, while self-supervised approaches leverage geometric and photometric consistency to adapt effectively to unseen environments [139].
Hybrid frameworks combining classical geometric models with deep neural networks offer a promising solution to challenges like scale ambiguity and drift. By integrating learned depth and optical flow predictions [140], these systems achieve consistent pose estimations, improving monocular system accuracy. For example, D3VO incorporates uncertainty estimation into the visual odometry pipeline for robust performance in dynamic and occluded environments. This highlights the value of merging data-driven and model-based methods for reliable localization [141, 142]. Recent advances in scene coordinate regression and neural implicit representations are transforming the field. Scene coordinate regression techniques eliminate the need for explicit 3D maps by learning direct transformations from image pixels to 3D coordinates, simplifying localization. Neural implicit representations, like NeRF [143] and Semantic-NeRF [144], encode scene geometry and semantics into compact neural fields, enabling robust localization in new environments and showcasing the potential of multimodal fusion for bridging perception and mapping. Despite progress, challenges remain, such as the need for large-scale datasets, high computational costs, and limited generalization to real-world conditions. Future research will likely focus on improving scalability, real-time capabilities, and developing unified frameworks that integrate localization with high-level perception and decision-making tasks [145].
@@ 2.5.1 Vision-Language-Action Models
Multimodal feature fusion is the core technology enabling the seamless transition from perception to action in robot manipulation tasks. Vision-Language-Action (VLA) models [159], which integrate visual perception, language understanding, and action planning, offer an efficient framework for addressing the challenges of complex manipulation tasks. The key innovation in VLA models lies in their ability to extract complementary features from different modalities and integrate them into a unified representation, enabling precise action planning and execution.
The visual modality is crucial for manipulation tasks, providing information about object position, shape, and layout. However, relying only on vision limits the robot’s task understanding. To address this, VLA models integrate language to process natural language instructions, bridging gaps in visual perception. For example, the RT-2 [160] model aligns visual and language representations via pretraining, allowing the robot to generate control commands from linguistic input using a Transformer architecture to map high-level semantics to actions. Additionally, the RoboMamba [161] model incorporates action dynamics into the multimodal fusion framework, optimizing vision, language, and action for task planning. The 3D-VLA [162] model enhances visual representation by generating 3D point clouds and depth maps, improving task precision and robustness with language-guided action generation in dynamic environments. OpenVLA [19] employs Low-Rank Adaptation (LoRA) to efficiently fine-tune VLMs, achieving cross-modal alignment with minimal trainable parameters. This design enables effective instruction following and generalization, particularly suited for deployment on resource-constrained robotic platforms. However, its focus on benchmark instruction tasks overlooks the continuous visual perception and fine-grained spatial reasoning needed in real-world robotics. Additionally, while LoRA ensures parameter efficiency, it may limit the model’s expressiveness in complex multimodal scenarios, such as indoor manipulation or dynamic path planning, where detailed spatial understanding is critical. VoxPoser [163] introduces a language-driven affordance reasoning framework that generates 3D value maps to guide complex manipulation. It excels in zero-shot generalization to novel objects and environments by linking natural language to 3D spatial actions. Yet, its reliance on voxel-based representations leads to high memory and compute costs, especially in large or high-resolution scenes. Moreover, the use of offline language models restricts real-time interaction and replanning, posing challenges for closed-loop control in dynamic robotic tasks.
@@ 3 Evolution of Vision-Language Models
The previous section provided a detailed overview of multimodal fusion strategies as applied to core robot vision tasks, such as semantic scene understanding, 3D object detection, SLAM, navigation, and manipulation. These fusion methods, grounded in traditional deep learning paradigms, lay the foundation for robust and efficient perception. However, recent advances in vision-language models (VLMs) introduce a new paradigm that enables robots to comprehend, reason, and act through natural language interfaces. In this section, we shift our focus to the evolution of vision-language models, examining their pretraining strategies, cross-modal alignment mechanisms, and emerging architectures that bridge visual and linguistic modalities for more generalizable robotic intelligence.
@@ 3.2 Cross-Modal Alignment and Representation Learning
Cross-modal alignment and representation learning are crucial for vision-language models, enabling different modalities (e.g., vision, text, audio) to interact in a shared embedding space. Alignment helps models map relationships between modalities, such as images to text or images to audio. Representation learning extracts useful features from each modality for various tasks. This combination improves performance in tasks like image captioning, visual question answering, and speech recognition, advancing multimodal capabilities. Cross-modal alignment maps information from modalities, such as images and text, into a shared space to understand their relationships [184, 185]. However, challenges arise from differences in structure, representation, and data features between modalities. Images are high-dimensional, while text consists of discrete symbols, making alignment difficult. Other challenges include scale differences and noise in large-scale unlabeled data, which can affect alignment accuracy [186].
Common methods for representation learning include contrastive learning, self-supervised learning, and cross-modal generative models, as well as emerging frameworks such as diffusion-based captioning models that provide new perspectives on vision-language generation [187]. Contrastive learning aims to minimize the distance between positive samples (same-category) and maximize it for negative samples (different-category). For cross-modal tasks, positive samples are visual and language pairs from the same context, while negative samples come from different contexts. The model is trained using a contrastive loss function to bring related modality pairs, like images and descriptions, closer in a shared embedding space while distancing unrelated pairs [188, 189]. The popular method CLIP [190] trains both vision and language models using contrastive learning, aligning images and text in a shared space for consistency. ALIGN [191] follows a similar approach, training cross-modal image and text data to enhance matching performance.
@@ 3.3 Transformer Variants and Large Vision-Language Models
In recent years, a new generation of VLMs has emerged, such as GPT-4V [216], BLIP-2 [223], and LLaVA-3D [224], demonstrating exceptional capabilities in open-vocabulary understanding, multimodal reasoning, and instruction following. However, their applicability in robot vision tasks remains constrained by several practical challenges. First, while these models excel in static image-text tasks, their limited generalization ability on robot-specific perceptual inputs (e.g., depth maps, 3D point clouds, and temporal data) stems from insufficient training during the pre-training phase. This limitation undermines their performance in tasks such as SLAM, dynamic scene understanding, and precise manipulation. Second, these models are typically large in size and rely on dense self-attention mechanisms, leading to high inference latency and memory consumption, making them difficult to deploy on resource-constrained platforms commonly used in robotics, such as embedded processors and edge devices. Additionally, these models often lack mechanisms for continuous adaptation in dynamic environments, a capability that is critical for long-term tasks such as autonomous navigation and interactive operations. To address these challenges, progress is needed in areas such as multimodal pre-training (covering 3D and temporal modalities), model compression (e.g., pruning, quantization, and adapter-based fine-tuning), and task-specific online learning strategies to achieve real-time, robust robot perception and decision-making capabilities.
@@ 4.1 Scene Understanding Datasets
In the domain of autonomous driving, nuScenes[23] and the Waymo Open Dataset[246] are two iconic datasets. nuScenes [23] offers data from six cameras, five radars, and one LiDAR, covering 360 views across 1,000 driving scenes, with more than 1.4 million frames annotated for 3D detection and tracking tasks across 23 categories. It captures diverse weather and lighting conditions in urban environments, but its limited geographic scope and relatively low vehicle density can affect the generalization performance in more complex or global traffic scenarios. The Waymo [246] dataset leads in terms of geographic coverage and data scale, containing 12 million 3D bounding box annotations and supporting cross-domain generalization and multimodal sensor fusion research. With synchronized high-resolution data from five LiDARs and five cameras, it supports comprehensive research in cross-domain generalization and multimodal sensor fusion. Its large volume, diverse locations, and rich annotations make it particularly suitable for real-world deployment and performance evaluation under varied driving conditions. Additionally, SemanticKITTI [24] focuses on semantic segmentation of LiDAR point clouds and provides an important benchmark for sequence-level semantic scene completion, with 43,000 densely annotated LiDAR scans. While it is valuable for understanding temporal consistency in point clouds, its reliance on a single sensor modality and limited environmental variability somewhat constrain its generalization to complex, sensor-rich real-world settings. The Cityscapes [247] dataset, with its high-resolution pixel-level annotations and complex urban street scene data, has become a crucial resource for urban semantic understanding. However, the dataset is limited to monocular RGB images captured from a fixed onboard camera, and lacks 3D information and sensor diversity, making it less suitable for tasks that require depth reasoning, multimodal fusion, or adaptation to dynamic urban scenes.
For indoor scene understanding, datasets such as Matterport3D[26], ScanNet [248], NYU Depth V2 [249], and Hypersim [250] have been pivotal. Matterport3D [26] provides 10,800 panoramic RGB-D views, along with accurate 3D reconstructions and semantic annotations, making it widely applicable for semantic segmentation, 3D reconstruction, and multi-task learning. ScanNet [248], through its 2.5 million RGB-D frames across 1,513 scenes, utilizes crowdsourced annotations to provide a large-scale, high-quality dataset for semantic segmentation and 3D modeling. NYU Depth V2 [249], as one of the earliest RGB-D datasets, offers 1,449 images from 464 diverse indoor scenes with detailed semantic segmentation and object relationship annotations, making it instrumental for research on physical support reasoning and cluttered environments. In addition, the Hypersim dataset [250], utilizing high-fidelity synthetic techniques, generates 77,400 indoor scene images with detailed pixel-level annotations, light decomposition, and surface normals, serving as an important tool for simulation-to-reality (Sim-to-Real) transfer research.However, despite offering valuable resources for visual perception tasks, these datasets still exhibit limitations in generalization to real-world robotic systems. Hypersim, being synthetic, lacks real sensor noise and physical interactions, leading to domain gaps in deployment. ScanNet and Matterport3D, while captured from real environments, are largely static and unimodal, making them less suitable for dynamic and interactive tasks. NYU Depth V2, though historically influential, has a relatively small scale and limited scene diversity, which constrains its applicability to modern multimodal and high-complexity scenarios.
@@
Datasets Scene Multimodal Data Data Scale Venue Year
360+x [251] Indoor, Outdoor Video/Audio 28 scenes, 1,688 videos CVPR 2024
ScanQA [252] Indoor RGB/Text 41,363 questions, 58,191 answers CVPR 2022
Hypersim [250] Indoor RGB/Depth 77,400 images ICCV 2021
NuScenes [23] Urban street RGB/Lidar/Radar 1.4 million frames annotated CVPR 2020
Waymo [246] Outdoor RGB/Lidar 12 million camera box annotations CVPR 2020
Semantickitti [24] Urban street RGB/Lidar 43,000 scans ICCV 2019
Matterport3D [26] Indoor RGB/Depth 10,800 panoramic RGB-D views arxiv 2017
ScanNet [248] Indoor RGB/Depth 1,513 scenes , 2.5 million RGB-D frames CVPR 2017
Cityscapes [247] Urban street 25,000 images RGB/Depth CVPR 2016
NYUDv2 [249] Indoor RGB/Depth 464 scenes , 1449 RGBD images ECCV 2012
@@ 4.2 Robot Manipulation Datasets
In the domain of multimodal perception, Touch and Go [255] and VisGel [256] have pioneered research into the integration of tactile and visual data. Touch and Go [255] collects paired visual and tactile data from real-world environments and explores cross-modal relationships using self-supervised learning methods. VisGel [256], leveraging GelSight sensors, constructs bidirectional generation tasks between tactile and visual modalities, further validating the feasibility of inter-modal information transformation. These datasets provide valuable experimental resources for enhancing multimodal perception research. However, compared to ObjectFolder 2.0 [257], these tactile datasets are relatively smaller in scale. ObjectFolder 2.0 [257], with its high-quality virtual object data and implicit neural representation technology, achieves efficient rendering of visual, auditory, and tactile modalities, supporting virtual-to-reality transfer learning tasks. This makes it a significant breakthrough in the field of multimodal perception learning.Although current multimodal robotic datasets increasingly emphasize generalization in their design, their practical adaptability to real-world scenarios still faces notable challenges. DROID [25] and RT-1 [254] are built on real-world data collection and exhibit strong task transferability and adaptability to natural environments, especially excelling in language-vision joint control. However, their scenarios are primarily confined to household and tabletop environments, with limited coverage of industrial, outdoor, or complex dynamic settings. Open X-Embodiment [260] constructs a large-scale cross-robot dataset by integrating 22 robotic platforms and over 500 tasks, marking a breakthrough in cross-platform transferability. Nevertheless, its task annotations remain coarse-grained, lacking precise modeling of fine-grained behaviors such as complex manipulation and semantic interaction, which may hinder its generalization in fine control tasks.In the grasping domain, R2SGrasp [253] leverages a sim-to-real framework to achieve effective real-world transfer, but its environmental backgrounds are limited and object diversity is constrained, restricting generalization to unseen scenarios. The Grasp-Anything series enhances semantic understanding through language modalities, improving generalization under complex instructions. Grasp-Anything++ [259], in particular, supports fine-grained tasks such as localized grasping and demonstrates semantic transfer capabilities under language grounding. However, its training still relies on synthetic data and a limited object set, leading to instability when encountering long-tail objects or ambiguous language commands in open-world settings. In the field of multimodal perception, Touch and Go [255] and VisGel [256] have expanded the frontier of vision-tactile integration. The former supports cross-modal self-supervised learning with real-world paired data, while the latter validates modality complementarity through bidirectional generation tasks between vision and touch. Nonetheless, both datasets are relatively small in scale and lack diversity across objects, environments, and interaction types, limiting their utility for training generalizable perception models. ObjectFolder 2.0 [257], although based on synthetic data, offers rich and high-quality multimodal representations, serving as a promising foundation for sim-to-real generalization. Still, how to reliably transfer such representations to the real world remains an open challenge.
In summary, these datasets have significantly advanced the generalization capabilities of robotic systems through multimodal fusion and expanded task coverage. However, notable gaps remain in data diversity, real-world complexity, and consistency in evaluation standards. Future research urgently requires unified benchmarks that encompass broader scenarios and behaviors, feature fine-grained annotations, and enable cross-task generalization assessments, thereby promoting the development of more universal and intelligent robotic systems.
@@ 5.1 Evaluation Metrics
where: - $mAP$ is the mean Average Precision, - $E_{i}$ represents the five error metrics: - mean Average Translation Error (mATE), - mean Average Scale Error (mASE), - mean Average Orientation Error (mAOE), - mean Average Velocity Error (mAVE), - mean Average Attribute Error (mAAE). These metrics ensure accurate assessment of multimodal fusion models’ capability to detect and localize objects.However, traditional AP-based metrics alone often fail to capture critical deployment-related factors such as temporal consistency and motion awareness. By integrating mAP with multiple error terms, NDS provides a more holistic evaluation that reflects not only detection accuracy but also localization robustness, size and orientation estimation precision, dynamic behavior modeling, and semantic attribute inference. This composite design aligns well with the real-world demands of autonomous driving and robotic perception systems, where safety, interpretability, and operational reliability are paramount.
@@
[…]
@@ 5.2 Performance comparison on 3D Object Detection
Methods Modality mAP mATE mASE mAOE mAVE NDS Car Truck C.V. Bus T.L. B.R. M.T. Bike Ped. T.C. VoxelNeXt [264] L 64.5 0.268 0.238 0.377 0.219 70.0 84.6 53.0 28.7 64.7 55.8 74.6 73.2 45.7 85.8 79.0 TransFusion-L [13] L 65.5 0.256 0.240 0.351 0.278 70.2 86.2 56.7 28.2 66.3 58.8 78.2 68.3 44.2 86.1 82.0 FocalFormer3D [265] L 68.7 0.254 0.242 0.340 0.218 72.6 87.2 57.0 34.4 69.6 64.9 77.8 76.2 49.6 88.2 82.3 UniM2AE-v2 [266] C 63.4 0.401 0.235 0.236 0.154 70.1 77.6 50.7 40.1 46.3 56.7 75.7 72.4 58.9 70.6 84.5 SparseBEV [267] C 60.3 0.425 0.239 0.311 0.172 67.5 76.3 49.2 35.6 44.2 53.4 76.8 64.0 53.7 65.8 83.7 RayDN [268] C 63.1 0.437 0.235 0.283 0.220 68.6 77.1 54.4 35.9 53.7 56.8 76.9 67.5 57.6 69.4 81.8 FusionPainting [87] C+L 68.1 0.256 0.236 0.346 0.274 71.6 87.1 60.8 30.0 68.5 61.7 71.8 74.7 53.5 88.3 85.0 AutoAlignV2 [269] C+L 68.4 0.245 0.233 0.311 0.258 72.4 87.0 59.0 33.1 69.3 59.3 78.0 72.9 52.1 87.6 85.1 IS-Fusion [94] C+L 76.5 0.230 0.231 0.267 0.223 77.4 89.8 67.8 44.5 77.6 68.3 81.8 85.3 65.6 93.4 91.1 UVTR-Multimodalit [88] C+L 67.1 0.306 0.245 0.351 0.225 71.1 87.5 56.0 33.8 67.5 59.5 73.0 73.4 54.8 86.3 79.6 3D-CVF [85] C+L 52.7 0.300 0.245 0.458 0.279 62.3 83.0 45.0 15.9 48.8 49.6 65.9 51.2 30.4 74.2 62.9 CMT [91] C+L 70.4 0.299 0.241 0.323 0.240 73.0 87.2 61.5 37.5 72.4 62.8 74.7 79.4 58.3 86.9 83.2 TransFusion [13] C+L 68.9 0.259 0.243 0.359 0.288 71.7 87.1 60.0 33.1 68.3 60.8 78.1 73.6 52.9 88.4 86.7 BEVFusion [92] C+L 70.2 0.261 0.239 0.329 0.260 72.9 88.6 60.1 39.3 69.8 63.8 80.0 74.1 51.0 89.2 86.5 DeepInteraction-e [90] C+L 75.6 0.235 0.233 0.328 0.226 76.3 88.3 64.3 44.7 74.2 66.0 83.5 85.4 66.4 92.5 90.9 SimpleBEV [270] C+L 75.7 0.236 0.235 0.283 0.143 77.6 89.0 69.2 75.2 70.5 44.4 91.7 81.2 69.1 87.9 78.4 DAL [271] C+L 72.0 0.253 0.238 0.334 0.174 74.8 89.1 60.2 73.3 65.8 34.6 89.6 81.7 58.5 86.6 80.6 SparseLIF [272] C+L 75.9 0.244 0.231 0.284 0.152 77.7 90.8 66.1 42.1 76.7 70.1 82.7 83.6 68.9 92.9 89.0 Ea-lss [273] C+L 76.6 0.234 0.228 0.278 0.204 77.6 90.2 67.1 43.9 76.7 69.1 84.1 85.9 66.6 91.3 91.2 MV2DFusion-e [274] C+L 77.9 0.237 0.225 0.247 0.192 78.8 91.1 69.7 45.7 76.7 70.4 83.0 87.1 72.2 93.3 90.3 RCBEVDet++ [275] C+R 67.3 0.341 0.234 0.241 0.147 72.7 83.5 56.0 45.0 54.1 57.9 76.9 78.4 60.2 76.4 85.0 CRN [105] C+R 57.5 0.416 0.264 0.456 0.365 62.4 78.4 52.6 29.2 57.0 48.3 64.7 63.4 47.8 58.2 74.9 RCM-Fusion [107] C+R 50.6 0.465 0.254 0.384 0.438 58.7 73.9 45.6 24.3 51.6 42.1 62.9 53.5 34.9 50.0 67.6 HyDRa [16] C+R 57.4 0.398 0.252 0.424 0.250 64.2 79.0 50.0 42.9 48.5 29.4 65.8 64.0 45.3 75.9 72.9 RaCFormer [276] C+R 64.9 0.358 0.240 0.329 0.179 70.2 83.0 56.0 39.2 53.6 56.8 78.3 70.6 55.6 72.0 83.9 CenterPoint v2 [277] C+L+R 67.1 0.249 0.236 0.350 0.250 71.4 87.0 57.3 28.8 69.3 60.4 71.0 71.3 49.0 90.4 86.8